Given that the mean and variance of a set of observations are 12.9 and 28.8528, determine the coefficient of variation
a.0.4471
b.0.4164
c.2.2367
d.2.4016
1
Expert's answer
2020-11-24T16:38:47-0500
"coefficient\ of\ variation=\dfrac{\sigma}{\bar{x}}="

"=\dfrac{\sqrt{28.8528}}{12.9}=0.4164"

b.0.4164